# General Compression Framework for Efficient Transformer Object Tracking

### v1

@@ 3.3 Prediction Guidance & Stage-wise Feature Mimicking
Replacement training enables the student model to learn the behavior of each individual stage, resulting in enhanced performance. However, merely forcing student model to emulate teacher model may be overly challenging for a smaller-sized student. Thus, we employ the teacher’s predictions to further guide the learning of compressed tracker. We apply the same loss as $L_{track}$ for prediction guidance, which is denoted as $L_{pred}$. With the aid of prediction guidance, student benefits from a quicker and stable learning process, assimilating knowledge from teacher model more effectively.
While prediction guidance accelerates the convergence, the student tracker might not entirely match the complex behavior of the teacher model. We introduce the stage-wise feature mimicking to further synchronize the temporal matching features between corresponding stages of the teacher and student models. This alignment is quantified by calculating the $L_{2}$ distance between the outputs of these stages, which is referred as $L_{feat}$. It is worth noting that any metric assessing the discrepancy in feature distributions can serve as the loss function. However, we choose a simple $L_{2}$ distance rather than a complex loss to highlight the effectiveness of our stage division and replacement training strategies. The stage-wise feature mimicking not only promotes a closer similarity in the feature representations of corresponding stages but also enhances the overall coherence between the teacher and student models.
Method LaSOT LaSOText TNL2K TrackingNet UAV123 FPS
AUC PNorm P AUC P AUC P AUC PNorm P AUC P
OSTrack-256 [49] 69.1 78.7 75.2 47.4 53.3 54.3 - 83.1 87.8 82.0 68.3 - 105
CompressTracker-2 60.4 87% 68.5 61.5 40.4 85% 43.8 48.5 89% 45.0 78.2 94% 83.3 74.8 62.5 92% 82.5 346 3.30×
CompressTracker-3 64.9 94% 74.0 68.4 44.6 94% 49.6 52.6 97% 50.9 81.6 98% 86.7 79.4 65.4 96% 88.3 267 2.54×
CompressTracker-4 66.1 96% 75.2 70.6 45.7 96% 50.8 53.6 99% 52.5 82.1 99% 87.6 80.1 67.4 99% 88.0 228 2.17×
CompressTracker-6 67.5 98% 77.5 72.4 46.7 99% 52.5 54.7 101% 54.3 82.9 99% 87.8 81.5 67.9 99% 88.7 162 1.54×
CompressTracker-8 68.4 99% 78.0 73.1 47.2 99% 53.1 55.2 102% 54.8 83.3 101% 88.0 81.9 68.2 99% 89.0 127 1.21×
Table 1: Compress OSTrack. We compress OSTrack multiple configurations with different layer settings. CompressTracker-x denotes the compressed student model with ’x’ layers. We report the performance on 5 benchmarks and calculate the performance gap in comparison to the original OSTrack. Our CompressTracker effectively achieves the balance between performance and speed.
Method LaSOT LaSOText TNL2K TrackingNet UAV123 FPS
AUC PNorm P AUC P AUC P AUC PNorm P AUC P
MixFormerV2-B [15] 70.6 80.8 76.2 50.6 56.9 57.4 58.4 83.4 88.1 81.6 69.9 92.1 165
MixFormerV2-S [15] 60.6 69.9 60.4 43.6 46.2 48.3 43.0 75.8 81.1 70.4 65.8 86.8 325
CompressTracker-M-S 62.0 88% 70.9 63.2 44.5 88% 47.1 50.2 87% 47.8 77.7 93% 82.5 73.0 66.9 96% 87.1 325 1.97×
Table 2: Compress MixFormerV2. We compress MixFormerV2 into CompressTracker-M-S with 4 layers, which is the same as MixFormerV2-S including the dimension of MLP layer. We report the performance on 5 benchmarks and calculate the performance gap in comparison to the origin MixFormerV2-B. Our CompressTracker-M-S outperforms MixFormerV2-S under the same setting.
@@ 3.4 Progressive Replacement
In Section 3.2, we describe the replacement training strategy. Although setting the Bernoulli sampling probability $p$ as a constant value can realize the compression, these stages have not been trained together at the same time and there may be some dissonance. A further finetuning step is necessary to achieve better harmony among the stages. Thus, we introduce a progressive replacement strategy to bridges the gap between the two initially separate training phases, fostering an end-to-end easy-to-hard learning process. By adjusting the value of $p$, we can control the number of stages to be replaced. The value of $p$ gradually increases from $p_{init}$ to $1.0$, allowing for a more incremental and coherent training progression:
$p=\left\{\begin{array}[]{ll}p_{init},&0<=t<\alpha_{1}m,\\ p_{init}+p_{init}\frac{t-\alpha_{1}m}{(1-\alpha_{1}-\alpha_{2})m},&\alpha_{1}m<=t<=(1-\alpha_{2})m,\\ 1.0,&(1-\alpha_{2})m<t<=m,\end{array}\right.$ (1)
@@ 3.5 Training and Inference
Our CompressTracker is a general framework applicable to a wide array of student model architectures. For the optimization of student model, our CompressTracker solely requires an end-to-end and easy-to-hand training process instead of multi-stage training methodologies. Furthermore, our approach simplifies the loss function design, eliminating the need for complex formulations. During training, teacher model is frozen and we only optimize student tracker. The total loss for CompressTracker is:
@@
Method LaSOT LaSOText TNL2K TrackingNet UAV123 FPS
AUC PNorm P AUC P AUC P AUC PNorm P AUC P
OSTrack-256 [49] 69.1 78.7 75.2 47.4 53.3 54.3 - 83.1 87.8 82.0 68.3 - 105
SMAT [22] 61.7 71.1 64.6 - - - - 78.6 84.2 75.6 64.3 83.9 158
CompressTracker-SMAT 62.8 91% 72.2 64.0 43.4 92% 46.0 49.6 91% 46.9 79.7 96% 85.0 75.4 65.9 96% 86.4 138 1.31×
Table 3: Compress OSTrack for SMAT. We compress OSTrack into CompressTracker-SAMT with 4 SMAT layers, which is the same as SMAT. We report the performance on 5 benchmarks and calculate the performance gap in comparison to the original OSTrack. Our CompressTracker-SAMT outperforms SMAT under the same setting.
@@ 4.1 Implement Details
Our framework CompressTracker is general and not dependent on a specific transformer structure, hence we select OSTrack [49] as baseline, which is a simple and effective transformer-based tracker. The training datasets consist of LaSOT [17], TrackingNet [35], GOT-10K [26], and COCO [32], following OSTrack [49] and MixFormerV2 [15]. We set $\lambda_{track}$ as $1$, $\lambda_{pred}$ as $1$, and $\lambda_{feat}$ as $0.2$. The $p_{init}$ is set as $0.5$. We train the CompressTracker with AdamW optimizer [34], with the weight decay as $10^{-4}$ and the initial learning rate of $4\times 10^{-5}$. The batch size is $128$. The total training epochs is 500 with 60K image pairs per epoch andthe learning rate is reduced by a factor of 10 after 400 epochs. $\alpha_{1}$ and $\alpha_{2}$ are set as 0.1. The search and template images are resized to resolutions of $288\times 288$ and $128\times 128$. We initialize the CompressTracker with the pretrained parameters of OSTrack. We report the inference speed on a NVIDIA RTX 2080Ti GPU.
Method LaSOT LaSOText TNL2K TrackingNet UAV123 FPS
AUC PNorm P AUC P AUC P AUC PNorm P AUC P
CompressTracker-2 60.4 68.5 61.5 40.4 43.8 48.5 45.0 78.2 83.3 74.8 62.5 82.5 346
CompressTracker-3 64.9 74.0 68.4 44.6 49.6 52.6 50.9 81.6 86.7 79.4 65.4 88.3 267
CompressTracker-4 66.1 75.2 70.6 45.7 50.8 53.6 52.5 82.1 87.6 80.1 67.4 88.0 228
CompressTracker-6 67.5 77.5 72.4 46.7 52.5 54.7 54.3 82.9 87.8 81.5 67.9 88.7 162
CompressTracker-8 68.4 78.0 73.1 47.2 53.1 55.2 54.8 83.3 88.0 81.9 68.2 89.0 127
HiT-Base [28] 64.6 73.3 68.1 44.1 - - - 80.0 84.4 77.3 65.6 - 175
HiT-Samll [28] 60.5 68.3 61.5 40.4 - - - 77.7 81.9 73.1 63.3 - 192
HiT-Tiny [28] 54.8 60.5 52.9 35.8 - - - 74.6 78.1 68.8 53.2 - 204
SMAT [22] 61.7 71.1 64.6 - - - - 78.6 84.2 75.6 64.3 83.9 158
MixFormerV2-S [15] 60.6 69.9 60.4 43.6 46.2 48.3 43.0 75.8 81.1 70.4 65.8 86.8 325
FEAR-L [6] 57.9 68.6 60.9 - - - - - - - - - -
FEAR-XS [6] 53.5 64.1 54.5 - - - - - - - - - 80
HCAT [10] 59.0 68.3 60.5 - - - - 76.6 82.6 72.9 63.6 - 195
E.T.Track [4] 59.1 - - - - - - 74.5 80.3 70.6 62.3 - 150
LightTrack-LargeA [45] 55.5 - 56.1 - - - - 73.6 78.8 70.0 - - -
LightTrack-Mobile [45] 53.8 - 53.7 - - - - 72.5 77.9 69.5 - - 120
STARK-Lightning [44] 58.6 69.0 57.9 - - - - - - - - - 200
DiMP [3] 56.9 65.0 56.7 - - - - 74.0 80.1 68.7 65.4 - 77
SiamFC++ [42] 54.4 62.3 54.7 - - - - 75.4 80.0 70.5 - - 90
Table 4: State-of-the-art comparison. We compare our CompressTracker which is compressed from OSTrack with previous light-weight tracking models. Our CompressTracker demonstrates superior performance over previous models.
@@ 4.2 Compress Object Tracker
In this section, we compress the pretrained OSTrack into different layer configurations. We report the performance of our CompressTracker across these configurations in Table 1. CompressTracker-4 compress OSTrack from 12 layers into 4 layers, and maintain $\mathbf{96\%}$ and $\mathbf{99\%}$ performance on LaSOT and TrackingNet while achieving $\mathbf{2.17\times}$ speed up. Furthermore, as shown in Figure 1, the training process of CompressTracker-4 is notably efficient, requiring only approximately 20 hours using 8 NVIDIA RTX 3090 GPUs. For CompressTracker-6 and CompressTracker-8, as we increase the number of layers, the performance gap between our compresstracker and OSTrack diminishes. It is worth noting that our CompressTracker even outperforms the origin OSTrack on some benchmarks. Specifically, CompressTracker-6 reaches $54.7\%$ AUC on TNL2K, and CompressTracker-8 achieves $55.2\%$ AUC on TNL2K and $83.3\%$ AUC on TrackingNet, while the origin OSTrack only achieves $54.3\%$ AUC on TNL2K and $83.1\%$ AUC on TrackingNet. Our framework CompressTracker demonstrates near lossless compression with the added benefit of increased processing speed.
Moreover, to affirm the generalization ability of our approach, we conduct experiments on MixFormerV2 [15] and SMAT [22]. MixFormerV2-S is a fully transformer tracking model consisting of 4 transformer layers, trained via a complex multi-stages model reduction paradigm. Following MixFormerV2-S, we adopt MixFormerV2-B as teacher and compress it to a student model with 4 layers. The results are shown in Table 2. Our CompressTracker-M-S share the same structure and channel dimension of MLP layers with MixFormerV2-S and outperforms MixFormerV2-S by about $1.4\%$ AUC on LaSOT. SMAT replace the vanilla attention in transformer layer with separated attention. We compress OSTrack into a student model CompressTracker-SMAT, aligning the number and structure of transformer layer with SAMT. We maintain the decoder of OSTrack for CompressTracker-SMAT. CompressTracker-SMAT surpasses SMAT by $1.1\%$ AUC on LaSOT, which demonstrates that our framework is flexible and not limited by the structure of transformer layer. Results in Table 123 verify the generalization ability and effectiveness of our framework.
4.3 Comparison with State-of-the-arts
To demonstrate the effectiveness of our CompressTracker, we compare our CompressTracker with state-of-the-art efficient trackers in 5 benchmarks. As shown in Table 4, our CompressTracker outperforms previous efficient trackers. Both HiT [28] and SMAT [22] are solely trained on the groundtruth and reduce computation through specialized network architectures. MixFormerV2-S [15] achieves model compression via a model reduction paradigm. Our CompressTracker-4 achieves $66.1\%$ AUC on LaSOT while maintaining 228 FPS. CompressTracker-4 outperforms HiT-Base by $1.5\%$ AUC on LaSOT without any specialized model structure design. CompressTracker-4 achieves the balance between speed and accuracy. Meanwhile, our CompressTracker-2, with just two transformer layers, maintains the highest speed at 346 FPS and also obtains competitive performance. CompressTracker-2 surpasses HiT-Tiny by $5.6\%$ AUC on LaSOT, and achieves about the same performance as MixFormerV2-S with only two transformer layers. As we add more transformer layers with CompressTracker-6 and CompressTracker-8, we see further improvements in performance. These outcomes demonstrate the effectiveness of our CompressTracker framework.
@@ 4.3 Comparison with State-of-the-arts
[…]
tableBackbone Initialization. ‘MAE-first4’ denotes initializing the student model using the first 4 layers of MAE-B. ‘OSTrack-skip4’ represents utilizing every fourth layer of OSTrack for the student model.
@@
[…]
tableDecoder Initialization and Optimization. ‘Random’ denotes randomly initialized decoder, and ’Teacher’ means the decoder is initialized with teacher parameters. ‘Frozen’ represents that the decoder is frozen, and ’Trainable’ denotes decoder is trainable.
@@
[…]
tableStage Division. ‘Even’ denotes evenly dividing stage strategy, and ‘Uneven’ means that the layer number of each stage in teacher model is 2,2,6,2.
@@
[…]
tableTraining Epochs. ’300’ and ’500’ denote the total training epochs.
@@ 4.4 Ablation Study
In this section, we conduct a series of ablation studies on LaSOT to explore the factors contributing to the effectiveness of our CompressTracker. Unless otherwise specified, the teacher model is OSTrack, and the student model has 4 encoder layers. The student model is trained for 300 epochs. Please see Appendix A.2 for more analysis.
@@
[…]
tableAblation studies on LaSOT to analyze the supervision of student model. The default choice for our model is colored in gray.
[…]
figureAblation study on training strategy.
Analysis on Supervision. We conduct a series of experiments to comprehensively analyze the supervision effects on the student model and to verify the effectiveness of our proposed training strategy. Results are presented in Table 4.4. Our proposed replacement training approach (#4) improves by 0.9 % AUC compared to singly training student model on groundtruth (#1), which demonstrates that the replacement training enhances the similarity between teacher and student models. Besides, prediction guidance (#5) and feature mimicking (#8) further boost the performance, indicating the effectiveness of the two strategies. Compared to only training on groundtruth (#1), our proposed replacement training, prediction guidance and feature mimicking collectively assist student model in more closely mimicking the teacher model, resulting in a total increase of $2.4\%$ AUC.
To further explore the generalization ability of our proposed training strategy, we compare the performance of models with different layer numbers and training settings, as illustrated in Figure 4.4. ’Naive Training’ denotes that the student model is trained without teacher supervision and replacement training. ’Distill Training’ represents that the student model is trained only with teacher supervision. ’CompressTracker’ refers to the same training setting in Table 4.4 #$\#8$. It can be observed that as the number of layers increases, there is a corresponding improvement in accuracy. Our CompressTracker shows a noticeable performance boost due to our proposed training strategy, which verifies the effectiveness and generalization ability of our framework.
@@
5 Limitation
While our CompressTracker demonstrates promising performance and generalization, its training is somewhat inefficient, requiring about 2×2$\times$2 × time compared to training a student model on ground truth data (20h vs. 8h on 8 NVIDIA 3090 GPUs, as shown in Figure 1 (a)). Moreover, a performance gap still exists between the teacher and student models suggests room for improvement in lossless compression. Future efforts will focus on developing more efficient training methods to boost student model accuracy and decrease training duration.
6 Broader Impacts
Our CompressTracker framework efficiently compresses object tracking models for edge device deployment but poses potential misuse risks, such as unauthorized surveillance. We recommend users to carefully consider the real-world implications and adopt risk mitigation strategies.

### v2

@@ 3.3 Prediction Guidance & Stage-wise Feature Mimicking
Replacement training enables the student model to learn the behavior of each individual stage, resulting in enhanced performance. However, mimicking the teacher directly can be too difficult for a smaller student model. Thus, we employ the teacher’s predictions to further guide the learning of compressed tracker. We apply the same loss as $L_{track}$ for prediction guidance, which is denoted as $L_{pred}$. This guidance accelerates and stabilizes the student’s learning, helping it assimilate knowledge from the teacher more effectively.
While prediction guidance accelerates the convergence, the student tracker might not entirely match the complex behavior of the teacher model. We introduce the stage-wise feature mimicking to further synchronize the temporal matching features between corresponding stages of the teacher and student models. This alignment is quantified by calculating the $L_{2}$ distance between the outputs of these stages, which is referred as $L_{feat}$. It is worth noting that any metric assessing the discrepancy in feature distributions can serve as the loss function. However, we choose a simple $L_{2}$ distance rather than a complex loss to highlight the effectiveness of our stage division and replacement training strategies. The stage-wise feature mimicking both promotes a closer similarity in feature representations of corresponding stages and enhances overall coherence between teacher and student.
@@ 3.4 Progressive Replacement
In Section 3.2, we describe the replacement training strategy. Although setting the Bernoulli sampling probability $p$ as a constant value can realize the compression, these stages have not been trained together at the same time and there may be some dissonance. A further finetuning step is necessary to achieve better harmony among the stages. Thus, we introduce a progressive replacement strategy to bridges the gap between the two initially separate training phases, fostering an end-to-end easy-to-hard learning process. By adjusting the value of $p$, we can control the number of stages to be replaced. The value of $p$ gradually increases from $p_{init}$ to $1.0$, allowing for a more incremental and coherent training:
$p=\left\{\begin{array}[]{ll}p_{init},&0\leq t<\alpha_{1}m,\\ p_{init}\frac{m+t-2\alpha_{1}m-\alpha_{2}m}{(1-\alpha_{1}-\alpha_{2})m}&\alpha_{1}m\leq t\leq(1-\alpha_{2})m,\\ 1.0,&(1-\alpha_{2})m<t\leq m,\end{array}\right.$ (5)
@@
Method LaSOT LaSOText TNL2K TrackingNet UAV123 FPS
AUC PNorm P AUC P AUC P AUC PNorm P AUC P
Model Generalization
CompressTracker-OSTrack [49] 66.1 96% 75.2 70.6 45.7 96% 50.8 53.6 99% 52.5 82.1 99% 87.6 80.1 67.4 99% 88.0 228 2.17×
CompressTracker-SeqTrack [12] 68.1 95% 77.3 73.4 47.9 96% 56.4 54.5 99% 54.2 83.1 98% 88.1 81.7 68.4 98% 90.1 78 1.95×
CompressTracker-ODTrack [52] 70.5 96% 79.5 76.9 50.9 97% 60.8 58.2 99% 58.9 82.8 97% 87.9 81.4 69.2 98% 90.8 87 2.71×
CompressTracker-SUTrack [13] 72.2 99% 80.7 78.2 52.1 98% 63.2 64.4 99% 69.9 84.3 99% 88.2 83.3 71.0 99% 91.4 134 2.42×
Stage Scalability
CompressTracker-OSTrack-2 60.4 87% 68.5 61.5 40.4 85% 43.8 48.5 89% 45.0 78.2 94% 83.3 74.8 62.5 92% 82.5 346 3.30×
CompressTracker-OSTrack-3 64.9 94% 74.0 68.4 44.6 94% 49.6 52.6 97% 50.9 81.6 98% 86.7 79.4 65.4 96% 88.3 267 2.54×
CompressTracker-OSTrack-4 66.1 96% 75.2 70.6 45.7 96% 50.8 53.6 99% 52.5 82.1 99% 87.6 80.1 67.4 99% 88.0 228 2.17×
CompressTracker-OSTrack-6 67.5 98% 77.5 72.4 46.7 99% 52.5 54.7 101% 54.3 82.9 99% 87.8 81.5 67.9 99% 88.7 162 1.54×
CompressTracker-OSTrack-8 68.4 99% 78.0 73.1 47.2 99% 53.1 55.2 102% 54.8 83.3 101% 88.0 81.9 68.2 99% 89.0 127 1.21×
Larger Transformer Scalability
CompressTracker-OSTrack-L-4 67.5 96% 77.9 72.7 45.9 98% 51.5 58.3 98% 58.5 83.2 99% 87.7 81.2 67.4 99% 88.2 228 2.84×
Higher Resolution Scalability
CompressTracker-OSTrack-384-4 67.7 96% 78.3 73.6 48.1 96% 58.0 54.3 99% 53.7 82.7 99% 87.3 81.0 68.2 98% 88.8 228 3.90×
Table 2: Generalization of CompressTracker. Our CompressTracker can be applied to any teacher model, any resolution, any stage number, any teacher model size, and any student architectures. CompressTracker-m-x denotes the compressed student model with ’x’ layers with ’m’ as teacher model. We report the performance on 5 benchmarks and calculate the performance gap in comparison to the origin teacher model.
3.5 Inner Connection of CompressTracker
The contributions of our CompressTracker are sequential and deeply interconnected for efficient transformer-based tracking. CompressTracker begins with the stage division strategy,is the foundation of our framework, as it directly enables replacement training, where student model stages are dynamically substituted with teacher model stages during training. Without this structured stage division, replacement training would not be feasible, underscoring the necessity of our initial design. Building on this, we further enhance the knowledge transfer process with prediction guidance and stage-wise feature mimicking. These components work in tandem, each reinforcing the next, ensuring that knowledge is transferred effectively at every stage. Our CompressTracker maintains a high level of efficiency throughout the compression process, which allows us to achieve an optimal balance between accuracy and efficiency, and highlights the inner connection and superiority of each contribution.
@@ 3.6 Training and Inference
Our CompressTracker is a general framework applicable to various student model architectures. It simplifies the student model optimization by using a straightforward end-to-end training process, without the need for multi-stage methods or complex loss functions. During training, teacher model is frozen and we only optimize student tracker. The total loss for CompressTracker is:
@@ 4.1 Implement Details
Our framework CompressTracker is general and not dependent on a specific transformer structure, hence we select a series of trackers, such as SUTrack [13], ODTrack [52], SeqTrack [12], OSTrack [49] and so on as baseline to verify the effectiveness of our CompressTracker. The training datasets consist of LaSOT [17], TrackingNet [37], GOT-10K [29], and COCO [34], following OSTrack [49] and MixFormerV2 [15]. We set $\lambda_{track}$ as $1$, $\lambda_{pred}$ as $1$, and $\lambda_{feat}$ as $0.2$. The $p_{init}$ is set as $0.5$. We train the CompressTracker with AdamW optimizer [36], with the weight decay as $10^{-4}$ and the initial learning rate of $4\times 10^{-5}$. The batch size is $128$. The total training epochs is 500 with 60K image pairs per epoch and the learning rate is reduced by a factor of 10 after 400 epochs. $\alpha_{1}$ and $\alpha_{2}$ are set as 0.1. The search and template images are resized to resolutions of $256\times 256$ and $128\times 128$. We initialize the CompressTracker with the pretrained parameters. We report the inference speed on a NVIDIA RTX 2080Ti GPU and Intel(R) Xeon(R) Platinum 8268 CPU @ 2.90GHz.
@@ 4.2 Compress Object Tracker
To validate the effectiveness of CompressTracker, we apply our CompressTracker framework to compress existing transformer-based tracking models, demonstrating its ability to balance performance and inference speed. We carefully select model configurations that offer an optimal trade-off between accuracy and efficiency. The experimental results are presented in Table 1. Our method achieves significant acceleration while preserving tracking performance. For instance, when compressing SUTrack (73.2 AUC on LaSOT), our CompressTracker maintains $99\%$ performance (72.2 AUC on LaSOT) while achieving a 2.42× speedup. Furthermore, compared to existing lightweight tracking models, CompressTracker consistently outperforms them in both speed and accuracy, highlighting its superiority and strength.
@@
Table 3: Backbone Initialization. ’MAE-first4’ denotes initializing the student model using the first 4 layers of MAE-B. ’OSTrack-skip4’ represents utilizing every fourth layer of OSTrack for the student model.
@@
Table 4: Decoder Initialization and Optimization. ’Random’ denotes randomly initialized decoder, and ’Teacher’ means the decoder is initialized with teacher parameters. ’Frozen’ represents that the decoder is frozen, and ’Trainable’ denotes decoder is trainable.
@@
Table 5: Stage Division. ’Even’ denotes evenly dividing stage, and ’Uneven’ means that the layer number of each stage in teacher model is 2,2,6,2.
# Replacement AUC Training Time
1 Random 65.2% 12 h
2 Decouple-300 64.6% 16 h
Table 6: Replacement training. ’Random’ denotes our replacement training, and ’Decouple-300’ represents decoupling the training of each stage.
# Replacement AUC
1 w/ Progressive 65.2%
2 w/o Progressive 64.8%
Table 7: Progressive Replacement. We conduct additional finetune if pregressive replacement is disabled.
@@
Table 8: Training Epochs. ’300’ and ’500’ denote total epochs.
# Model Training Time
1 CompressTracker-4 20 h
2 OSTrack 17 h
3 MixFormerV2-S 120 h
Table 9: Training Time comparison with other methods.
# Method AUC FPS
1 CompressTracker-4 66.1 228
2 Distillation 63.8 228
3 Pruning (MixFormerV2-S) 60.6 325
Table 10: Comparison with Other Compression Techniques.
@@ 3 MixFormerV2-S 120 h
4.3 Generalization Verification
In the previous section, we demonstrate the effectiveness of our CompressTracker. To further validate its generalization ability, we conduct additional experiments across diverse settings. The results, shown in Table 2, confirm that our method maintains strong performance across different model architectures, stage numbers, input resolutions, and parameter scales. We also show the results with different structure of student in Supplementary Materials. Moreover, CompressTracker is highly efficient in training, requiring only 20 hours for CompressTracker-OSTrack-4 using 8 NVIDIA RTX 3090 GPUs, in contrast to the 120 hours needed for MixFormerV2-S. This training efficiency further highlights the strong generalization capability of our method, as it can rapidly adapt to various configurations while maintaining high accuracy. Based on the extensive experiment results, our CompressTracker has both highly effectiveness and stronge generalization. CompressTracker maximizes performance retention while significantly improving inference speed and is applicable to any teacher model, any resolution, any stage number, any teacher model size, and any student architectures, which previous methods cannot achieve.
@@ 4.4 Ablation Study
In this section, we conduct a series of ablation studies on LaSOT to explore the factors contributing to the effectiveness of our CompressTracker. Unless otherwise specified, the teacher model is OSTrack, and the student model has 4 encoder layers. The student model is trained for 300 epochs and $p_{init}$ is set as 0.5. Please see Supplementary Materials for more analysis.
Figure 3: Ablation study on different replacement probability.
Figure 4: Ablation study on training strategy.
@@
Table 12: Ablation study on LaSOT about supervision of student model. The default choice for our model is colored in gray.
Analysis on Supervision. We conduct a series of experiments to comprehensively analyze the supervision effects on the student model and to verify the effectiveness of our proposed training strategy. Results are presented in Table 12. Our proposed replacement training approach (#4) improves by 0.9 % AUC compared to singly training student model on groundtruth (#1), which demonstrates that the replacement training enhances the similarity between teacher and student models. Besides, prediction guidance (#5) and feature mimicking (#8) further boost the performance, indicating the effectiveness of the two strategies. Compared to only training on groundtruth (#1), our proposed replacement training, prediction guidance and feature mimicking collectively assist student model in more closely mimicking the teacher model, resulting in a total increase of $2.4\%$ AUC.
To further explore the generalization ability of our proposed training strategy, we compare the performance of models with different layer numbers and training settings, as illustrated in Figure 4. ’Naive Training’ denotes that the student model is trained without teacher supervision and replacement training. ’Distill Training’ represents that the student model is trained only with teacher supervision. ’CompressTracker’ refers to the same training setting in Table 12 #$\#8$. It can be observed that as the number of layers increases, there is a corresponding improvement in accuracy. Our CompressTracker shows a noticeable performance boost due to our proposed training strategy, which verifies the effectiveness and generalization ability of our framework.
Replacement Training. To evaluate the efficiency and effectiveness of our replacement training strategy, we conduct experiments presented in Table 6. ’Random’ denotes our replacement training, and ’Decouple-300’ represents stage-by-stage decoupling. Result of # 1 aligns with our replacement training with 300 training epochs, while in # 2, we apply decoupled training, sequentially training and freezing each stage for 75 epochs, followed by 30 epochs of fine-tuning. The ’Decouple-300’ (# 2) approach achieves $64.6\%$ AUC on LaSOT with the same training epochs, marginally lower by $0.6\%$ AUC than our replacement training strategy (# 1). The ’Decouple-300’ approach (# 2) requires a complex, multi-stage training along with supplementary fine-tuning, which may suffer from suboptimal outcomes at a specific training process. However, our CompressTracker operates on an end-to-end, single-step basis, and can avoid the suboptimal performance issue through its unified training manner, which validates the superiority of our replacement training.
@@
Training Time. We compare the training time of our CompressTracker-4 with 500 training epochs, OSTrack, and MixFormerV2-S in Table 11. The training time is recorded on 8 NVIDIA RTX 3090 GPUs. Although our CompressTracker requires a longer training time compared to the OSTrack, the increased computational overhead remains within acceptable limits. Moreover, MixFormerV2-S is trained on 8 Nvidia RTX8000 GPUs, and we estimate this will take roughly 80 hours on 8 NVIDIA RTX 3090 GPUs based on the relative computational capabilities of these GPUs. The training time of our CompressTracker-4 is significantly less than that of MixFormerV2-S, which validate the efficiency and effectiveness of our framework.